What is the molecular mass (Formula mass) if a 3.2 g of the gas occupies a volume of 1.6 L at 30 degree Celsius and 745 Millimeter mercury?


1
Expert's answer
2021-01-28T05:09:39-0500
"P1V1\/T1=P0V0\/T0"

745 mm Hg = 99.325 kPa


"V0=P1V1T0\/(T1*P0)"

"V0=99.325*1.6*273\/(303*101.325)=1.4 L"

"V\/Vm=m\/M"

"M=3.2*22.4\/1.4=51.2 g\/mol"